OpenStudy (anonymous):

A car traveling at 60 km/h will skid 30 m when its brakes are locked. If the same car is traveling at 180 km/h, what will be its skidding distance? Show your work.

OpenStudy (aaronq):

i guess, if you assume that it's proportional, then you can build a ratio: \(\dfrac{60\;km/h}{30\;m}=\dfrac{180\;km/h}{x\;m}\)

OpenStudy (anonymous):

When I solve that am I supposed to see how many times 30 goes into 60 and 180?

OpenStudy (anonymous):

\[\frac{ 2km/h }{ m } = \frac{ 6km/h }{ x m }\] ?

OpenStudy (anonymous):

Yes, that works, but I would convert meters to kilometers, or vice versa just to make the calculation easier.

OpenStudy (anonymous):

Are 500 km equivalent to 1 m?

OpenStudy (anonymous):

1km = 1000m

OpenStudy (anonymous):

\[\frac{ 2km/h }{ 1000km } = \frac{ 6km/h }{ 1000km \times x }\] ?

OpenStudy (anonymous):

No, you converted made 1 m = 1000m; let us start from the beginning. Write your original first equation down.

OpenStudy (anonymous):

\[\frac{ 60km/h }{ 30m } = \frac{ 180km/h }{ x m }\]

OpenStudy (anonymous):

For simplicity's sake, we shall convert km to m. 1 km is 1000m. So 60 and 180 become 60,000 and 180,000 respectively.

OpenStudy (anonymous):

\[\frac{ 60,000m/h }{ 30m } = \frac{ 180,000m/h }{ x m }\]

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

\[\frac{ 20,000m/h }{ m } = \frac{ 60,000m/h }{ xm }\]

OpenStudy (anonymous):

Keep the 180,000

OpenStudy (anonymous):

\[\frac{ 20,000m/h }{ m } = \frac{ 180,000m/h }{ xm }\]

OpenStudy (anonymous):

good, now cross multiply, keep the units, u should see the unit "m^2/h"

OpenStudy (anonymous):

\[\frac{ 20,000m ^{2}/hx }{ 60,000m^2/h }\] I'm not confident with the placement of the x value

OpenStudy (anonymous):

Let's go back to your previous step. 20,000 xm^2/h = 60,000 m^2/h is what you get after cross multiplying.

OpenStudy (anonymous):

Oooooh my fault... I see

OpenStudy (anonymous):

\[\frac{ 20,000m ^{2}/hx }{ 60,000m^2/h }\] I'm not confident with the placement of the x value

OpenStudy (anonymous):

Oooooh my fault... I see

OpenStudy (anonymous):

now divide each side by 20,000 m^2/h (from my equation)

OpenStudy (anonymous):

Would that be 30,000m/hx

OpenStudy (anonymous):

AH, I know where you went wrong. 60,000/30 is not equal to 20,000/1

OpenStudy (anonymous):

So, 2,000 xm^2/h = 60,000 m^2/h is what you actually get after cross multiplying.

OpenStudy (anonymous):

make that 60,000 a 180,000 actually

OpenStudy (anonymous):

90m, is your final answer.

OpenStudy (anonymous):

Thank you so much for going step by step with me ♥

OpenStudy (anonymous):

Glad to well :).

OpenStudy (anonymous):

help*

OpenStudy (anonymous):

If you could I'd appreciate some help on a few more questions but I understand if you can't, thank you again. [:

OpenStudy (anonymous):

ask away, I'll see what I can do

OpenStudy (anonymous):

thank you